OpenStudy (amandahhhhhhh):

Last year, Lamar had $30,000 to invest. He invested some of it in an account that paid 6% simple interest per year, and he invested the rest in an account that paid 8% simple interest per year. After one year, he received a total of $2020 in interest. How much did he invest in each account?

OpenStudy (amandahhhhhhh):

please help me

OpenStudy (3mar):

Salam First: the 30,000 is divided into two amounts; x and y. so its equation will be : \[x+y=30,000\] Second: the rate of 1st amount is 0.06x and the for the other amount is 0.08y, so the total amount of the invest is expressed in the form : \[0.06x + 0.08y = 2020\] So you can easily solve these two equations together \[x+y=30,000\] \[0.06x + 0.08y = 2020\] You will get that : \[x=19,000\] \[y=11,000\] To be sure, add x to y, it will lead to 30,000 Done!